At some point in the life of a project, the project manager determined the following data on a $1,250,000 authorized budget project: amount of earned value $350,000. At that point, the value of the planned work was $750,000, and actual cost was $750,000.

Using the data provided above, and assuming a linear cost function, the project manager can estimate that actual cost of the project would be closest to (figures rounded to nearest hundred thousand):

  1. $4,500,000
  2. $2,700,000
  3. $2,400,000
  4. $2,100,000
Answer #1

Cost Variance = Actual cost - Planned cost = 750000-750000= 0

Performance Index of project =Earned Value/Planned Value = 350000/750000 = 0.47

actual cost of the project = $1,250,000 / 0.47 = $ 2,659,574

which close to option b:$2,700,000